# SuperCarver: Texture-Consistent 3D Geometry Super-Resolution for High-Fidelity Surface Detail Generation

### v1

@@ 3.1 Overall Workflow
Given a 3D textured model $\mathcal{M}$ with vertices $\mathcal{V}$ and faces $\mathcal{F}$, we aim at generating high-fidelity surface details to produce the updated geometric mesh structure $\mathcal{M}^{u}=(\mathcal{V}^{u},\mathcal{F}^{u})$. As illustrated in Figure 3, we start by rendering $\mathcal{M}$ from certain specified viewpoint to obtain the corresponding color image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, normal map $\mathcal{N}\in\mathbb{R}^{H\times W\times 3}$, and depth map $\mathcal{D}\in\mathbb{R}^{H\times W}$. We achieve geometry super-resolution with a two-stage pipeline. In the first stage, we construct a deterministic prior-guided 2D diffusion architecture to predict a high-precision normal map $\mathcal{H}\in\mathbb{R}^{H\times W\times 3}$. By performing inferences from multiple different viewpoints, we can acquire a set of multi-view normal map predictions $\{\mathcal{H}_{k}\}_{k=1}^{K}$. In the second stage, we design a noise-resistant inverse normal rendering approach to iteratively optimize the original mesh surface with geometric constraints, where we perform distance field deformation by learning grid offsets. In what follows, we will detail the two stages.
3.2 Deterministic Prior-Guided Normal Diffusion
We resort to the most popular text-to-image latent diffusion architecture [87] to leverage its strong generalization power and adapt it to generate texture-consistent detail-rich normal maps. Some previous studies [58, 91, 59, 113] resort to joint multi-view diffusion to mitigate inconsistency across views. Still, such a modeling paradigm also restricts the flexibility in adjusting the number and positions of viewpoints if users have personalized needs to designate specific surface areas for geometric detail carving, which are common in practical scenarios. Moreover, since in our setting there is no need to infer invisible novel views from limited visual observations, inter-view inconsistency will not be a major problem. Thus, we choose the more flexible single-view working mode.
@@ 3.2 Deterministic Prior-Guided Normal Diffusion
To achieve this goal, the most straightforward solution is to denoise the noisy latent representation of the target normal map $\mathcal{H}$ while incorporating the color image $\mathcal{I}$ into the denoising U-Net. Thus, the forward process is given by:
@@
where the noise schedule term $\bar{\alpha}_{t}$ controls the transformation to the standard Gaussian distribution $\epsilon$, $T$ is the number of time steps, $\widetilde{\mathcal{H}}$ denotes the latent representation of $\mathcal{H}$ after VAE’s encoder, and ~$\widetilde{\mathcal{H}}_{t}$ is the noisy feature map.
@@
where $\mathbf{c}$ is a vectorized embedding originally extracted from texts. The visual cues can be injected via concatenating ~$\widetilde{\mathcal{H}}_{t}$ with the encoded latent feature map (denoted as $\widetilde{\mathcal{I}}$) of $\mathcal{I}$.
Deterministic Diffusion. In fact, the above implementation together with various existing approaches [59, 27, 113] can be regarded as cross-domain (RGB2Normal) image translation, while our special setting motivates us to directly refine the rendered coarse normal map $\mathcal{N}$ to generate the desired high-precision normal map $\mathcal{H}$. To this end, we investigate a deterministic diffusion mechanism, whose forward process can be formulated as:
@@
where we directly replace the Gaussian noise $\epsilon$ (in Eq. (1)) with the latent representation $\widetilde{\mathcal{N}}$ of the rendered normal map $\mathcal{N}$ after VAE’s encoder.
Empirically, eliminating the inherent stochastic nature in this way naturally ensures deterministic outputs and avoids the fluctuations of diffused normal maps across views, thus implicitly enhancing inter-view consistency and stabilizing the subsequent multi-view inverse rendering process.
Prior Extraction. To boost the accuracy and robustness of normal diffusion, we further extract semantic and geometric priors as conditioning signals.
For the rendered color image $\mathcal{I}$, we extract its DINO [75] feature $\widetilde{\mathcal{Q}}$ and also obtain its latent representation $\widetilde{\mathcal{I}}$ through VAE’s encoder. The spatial resolution of $\widetilde{\mathcal{Q}}$ is pre-upscaled by several auxiliary convolutional layers. Besides, we apply CLIP’s image encoder to extract the vectorized embedding $\tilde{\mathbf{c}}$ for cross-attention within U-Net’s transformer blocks. For the rendered depth map $\mathcal{D}$, we treat its latent representation $\widetilde{\mathcal{D}}$ as the geometric prior. Then we concatenate $\widetilde{\mathcal{Q}}$, $\widetilde{\mathcal{I}}$, and $\widetilde{\mathcal{D}}$ to deduce the resulting conditioning signal $\widetilde{\mathcal{S}}$.
@@
Empirically, we introduce another parameterization form of v-prediction [88] to train the denoising U-Net, whose input channels and cross-attention dimensions are adapted to match with our prior-conditioned latent representations and CLIP-encoded image embeddings.
3.3 Noise-Resistant Inverse Normal Rendering
@@
When finishing the training of our normal diffusion model, we can render multi-view images from different viewpoints $\{\mathbf{v}_{k}\}_{k=1}^{K}$, and perform inference to correspondingly predict a set of detail-rich normal maps $\{\mathcal{H}_{k}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{K}$.
@@
To transfer the underlying 3D structural patterns implied by the predicted 2D normal maps to the input mesh surface, the common practice is to perform gradient-based mesh refinement through inverse rendering. However, it is observed that updating the mesh geometry from potentially imperfect normal map predictions is prone to causing obvious surface degradation, even with adaptations of various existing state-of-the-art (explicit [76] or implicit [90]) mesh optimization approaches. Hence, we are strongly motivated to develop a highly noise-resistant scheme better suited to our setting of adding new details to a given coarse surface.
@@
We introduce a simple yet effective distance field deformation approach, working in a “value-fixing, grid-moving” manner. Specifically, assuming that the original mesh $\mathcal{M}=(\mathcal{V},\mathcal{F})$ has already been normalized into the unit sphere, we start by constructing uniform 3D grid points $\mathcal{G}\in\mathbb{R}^{r\times r\times r\times 3}$ within $[-1,1]^{3}$ and computing per-grid signed distance values $\mathcal{X}\in\mathbb{R}^{r\times r\times r}$. Different from common ways of implementing deformable distance field representations via only updating $\mathcal{X}$ or jointly updating $\mathcal{G}$ and $\mathcal{X}$, we choose a subtle strategy that fixes distance values $\mathcal{X}$ while separately learning zero-initialized offsets $\mathcal{O}\in\mathbb{R}^{r\times r\times r\times 3}$ to move initial grid positions $\mathcal{G}$. In comparison, standard distance field refinement requires higher resolutions of fixed uniform grids; otherwise, detailed structures cannot be captured even with accurate distance values. Additionally, fixing pre-computed distance values hinders the optimization process from drastically changing the original mesh structure even with obviously erroneous normal prediction in certain areas.
@@
To achieve this, we employ a differentiable implementation of a classic iso-surface extraction algorithm DMC [74], which we denote as $\Psi(\cdot;\cdot)$, for producing an updated mesh $\mathcal{M}^{u}=(\mathcal{V}^{u},\mathcal{F}^{u})$. Then, a differentiable rasterizer $\mathcal{R}(\cdot;\cdot)$ is integrated to render the correspondingly updated normal maps $\{\mathcal{H}_{k}^{u}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{K}$ w.r.t. every viewpoint $\mathbf{v}_{k}$. The whole processing procedure can be formulated as:
@@
where the first term aligns every pair of diffusion-predicted and mesh-rendered normal maps, the last two terms denote mesh vertex and normal regularizations weighted by factors $\omega_{s}$ and $\omega_{n}$. For each vertex $\mathbf{p}_{i}$ in the updated mesh and its neighboring vertices $\Omega_{i}=\{\mathbf{p}_{i,j}\}$, the Laplacian smoothing is computed by:
@@
After iterative optimization, we moderately apply Taubin smoothing [102] as post-processing to mitigate meaningless surface wrinkles.
4 Experiments
Figure 5: Visual comparison of normal map prediction results exported from (c) GeoWizard [27], (d) StableNormal [119], and (e) Ours, where (a) and (b) show the corresponding color images and ground-truth high-fidelity normal maps.
Table 1: Quantitative comparison of normal map prediction accuracy under the metric of mean angular error (MAE).
@@ 4 Experiments
Figure 6: The rendering effects of (c) our optimized mesh and (d) attaching the bump map obtained from PBRBoost [76], given the input textured coarse mesh as presented in (a) and (b).
Training Dataset Preparation. We used the publicly available Objaverse [23] dataset comprising 800K+ 3D assets to produce our training data. Through a series of meticulously designed filtering protocols and cleaning procedures, we ultimately acquired around 80K high-quality mesh models.
To create paired low-poly and high-poly normal renderings, we performed geometry simplification on each high-poly training mesh model based on manifold reconstruction [36], and further removed detailed structures using [102]. We also manually annotated a smaller portion of unqualified training samples, which were used for training a separate neural model to further purify the whole training set.
Implementation Details. We adopted the popular diffusion architecture of Stable Diffusion V2.1 [87] and performed fine-tuning on 16 NVIDIA Tesla V100 GPUs for 40K steps with a total batch size of 128. We applied the AdamW [63] optimizer with cosine annealing, reducing the learning rate from 5e-5 to 1e-5. For both training and inference, the image resolution is configured to 512$\times$512.
For iterative mesh sculpting, we adopted Nvdiffrast [43] for efficient differentiable rendering from 12 camera poses (8 evenly-spaced horizontal views with azimuth angles from 0 to 360 and 4 downward-looking (elevated by 30) views with azimuth angles [45, 135, 225, 315]). Camera intrinsics were kept the same as our configurations in BlenderProc. The hyperparameters were empirically tuned as $r=512$, $\tau=0.5$, $\omega_{s}=0.25$, $\omega_{n}=0.01$ for 200 iterations.
4.1 Performance Evaluation
Table 2: Normal diffusion accuracy of different baseline schemes. Stoc. denotes the straightforward implementation of stochastic diffusion. w/o Sem. and w/o Geo. respectively denote removing the high-level semantics and low-level depth priors.
@@ 4.1 Performance Evaluation
Considering the lack of specialized research on our targeted task of texture-consistent 3D geometry super-resolution, we evaluated our superiority via applying SuperCarver on the textured meshes generated by various existing state-of-the-art 3D generation approaches, including InstantMesh [117], Unique3D [113], and CLAY [127], covering representative generation paradigms of feed-forward large reconstruction models, multi-view diffusion models, and native 3D latent diffusion models. As demonstrated in Figure 4, various input meshes are effectively boosted by our approach to produce realistic and expressive 3D surface details well aligned with the corresponding texture appearances.
Comparison on Normal Prediction. To evaluate the effectiveness of our specialized normal diffusion architecture, we made comparisons with advanced normal prediction frameworks, including GeoWizard [27] and StableNormal [119]. To facilitate quantitative evaluations, we manually selected 50 detail-rich textured meshes as ground-truths from Objaverse [23] and ABO [21] datasets. Then we applied surface simplification and smoothing to create geometrically-coarse meshes as the inputs.
As shown in Figure 5, our normal map predictions show better details and avoid obviously erroneous areas. Besides, we further deduced the metric of mean angular error (MAE) between ground-truth and predicted normal maps, where we masked invalid backgrounds for the outputs of GeoWizard and StableNormal. As illustrated in Table 1, owing to our specialized diffusion architecture, our normal map predictions significantly outperform the competing approaches. In addition, we also considered PBRBoost [107] (a latest material generation framework), since one of its outputs is the 2D bump maps associated with the base meshes, enabling us to compare its rendering effects when attaching the bump maps as height information. As shown in Figure 6, the actual fidelity of the enhanced geometric cues from PBRBoost is still far from satisfactory.
Comparison on Mesh Optimization. To facilitate surface geometry updating, many recent 3D generation frameworks [64, 20, 50, 113] introduce the CR [76] algorithm with joint optimization and remeshing, which is regarded as the most relevant competitor. In fact, we had previously attempted to adapt FlexiCubes [90], an advanced iso-surface representation, to achieve gradient-based geometry refinement starting from the initially given mesh surface. However, through extensive attempts, we empirically found that it easily leads to severe surface damages. We reason that this is attributable to its significantly increased deformation flexibility (which should have been its major strength) amplifies the impact of erroneous normal pixels. As shown in Figure 7, for typical hard cases with less accurate normal predictions, CR tends to produce noisy and degraded surfaces.
4.2 Ablation Study
To evaluate the effectiveness and necessity of the core components in our entire workflow, we conducted ablation study with different baseline schemes. We began with a straightforward baseline normal diffusion framework as formulated in Eq. (1) and Eq. (2). Besides, since prior guidance is also a critical condition for generating robust and accurate normal maps, we can design another two variants via removing the semantic cues extracted from DINO features or the geometric cues extracted from depth maps. As reported in Table 2, directly fine-tuning a standard diffusion model conditioned on color images lead to significantly degraded performance, with the MAE metric of 21.64. More importantly, as illustrated in Figure 8, the stochastic nature inevitably results in unstable predictions. As reported in the last two columns of Table 2, the joint injection of semantic and geometric priors are indeed beneficial. In the inverse normal rendering stage, we experimented with another two distance field deformation schemes, i.e., fixing grids while optimizing distances; jointly updating grids and distances. To obtain quantitative results, we computed the MAE metric of normal maps rendered from optimized meshes (note that Table 1 and Table 2 are targeted at diffused normal maps). As evidenced in Table 3, the two variants show sub-optimal performances, in that only optimizing distances usually causes weakened details while the joint optimization strategy easily suffers from surface distortion especially when the diffused normal maps contain relatively larger errors.
Table 3: Quantitative comparison of different distance field deformation schemes. Only Grids means our original design, Only Dist. and Joint G&D are the corresponding two baseline variants.
@@ 4.2 Ablation Study
5 Conclusion

### v2

@@ III-A Overall Workflow
Figure 2 illustrates the overall processing procedures of our proposed SuperCarver. Technically, we begin by rendering the input textured mesh $\mathcal{M}$ from a certain pre-specified viewpoint, respectively producing the corresponding 2D RGB color image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, normal map $\mathcal{N}\in\mathbb{R}^{H\times W\times 3}$, and depth map $\mathcal{D}\in\mathbb{R}^{H\times W}$. The subsequent 3D geometry super-resolution is achieved in two sequential stages. In the first stage, we propose a deterministic prior-guided diffusion architecture conditioned on diverse semantic and geometric cues to predict a detail-rich normal map $\mathcal{H}\in\mathbb{R}^{H\times W\times 3}$. Thus, by performing inferences from multiple different viewpoints, we can acquire a group of multi-view normal map predictions, denoted as $\{\mathcal{H}_{k}\}_{k=1}^{K}$. In the second stage, we propose a noise-resistant inverse normal rendering approach based on deformable distance field representation to iteratively optimize mesh surface with geometric constraints. In what follows, we will detail the two stages and specifically explain our design choices.
III-B Deterministic Prior-Guided Normal Diffusion
We resort to the most popular text-to-image latent diffusion architecture [5] for leveraging its strong generalization power and adapting it to generate texture-consistent detail-rich normal maps. Some previous studies [10, 48, 11, 12] resort to joint multi-view diffusion to mitigate inconsistencies across views, yet in our implementation we choose the single-view working mode based on the following considerations:
  • 1)
    Our targeted task of normal boosting estimates the detail patterns reflected by texture appearances based on given coarse normal maps, making the actual learning process more “predictive” than purely “generative”. Moreover, there is no need to infer invisible contents and generate novel views from limited visual observations. Therefore, inter-view inconsistency can be expected to be controlled within reasonable ranges.
  • 2)
    The joint multi-view learning architecture requires fixed view configuration pre-determined during training. This setup restricts the flexibility in adjusting the number and positions of viewpoints if users have personalized needs (which turn to be quite common in practice) to designate specific surface areas for detail sculpting.
@@ III-B Deterministic Prior-Guided Normal Diffusion
Straightforward Solution. To achieve this goal, a straightforward diffusion-based solution is to denoise the noisy latent representation of the target normal map $\mathcal{H}$ while incorporating the color image $\mathcal{I}$ into the denoising U-Net. Thus, the forward process can be formulated as:
@@
where the noise schedule term $\bar{\alpha}_{t}$ controls the transformations to the standard Gaussian distribution $\epsilon$, $T$ is the number of time steps, $\widetilde{\mathcal{H}}$ denotes the latent representation of $\mathcal{H}$ after VAE’s encoder, and ~$\widetilde{\mathcal{H}}_{t}$ is the noisy feature map.
@@
where $\mathbf{c}$ is the vectorized embedding originally extracted from textual condition, and the visual cues can be injected by means of concatenating ~$\widetilde{\mathcal{H}}_{t}$ with $\widetilde{\mathcal{I}}$, the encoded 2D latent feature map of the color image $\mathcal{I}$.
Deterministic Diffusion. In fact, such straightforward implementation as introduced above, together with various existing generative geometry modeling approaches [11, 24, 12], can be regarded as cross-domain (RGB2Normal) image-to-image translation, while the special task of geometry super-resolution motivates us to refine the rendered coarse normal map $\mathcal{N}$ to generate the desired high-precision normal map $\mathcal{H}$.
To achieve stable and accurate normal detail generation, we introduce a deterministic diffusion mechanism by shifting the source and target domains, modeling a sequence of interpolations between coarse and target normal maps. Intuitively, this learning process can also be understood as multi-step image-to-image translations. In this way, the corresponding forward process can be formulated as:
@@
where we directly replace the Gaussian noise $\epsilon$ (as in Eq. (1)) with the latent representation $\widetilde{\mathcal{N}}$ of the rendered normal map $\mathcal{N}$ embedded from the VAE’s encoder.
Eliminating the inherent stochastic nature of diffusion naturally ensures deterministic outputs and avoids the fluctuations of the generated normal maps when inferenced from multiple views, implicitly enhancing inter-view consistency and stabilizing the subsequent multi-view inverse rendering process.
Prior Extractions. To enhance the accuracy and robustness of normal diffusion, we further extract semantic and geometric priors serving as the conditioning signals.
For the rendered color image $\mathcal{I}$, we extract its DINO [85] feature $\widetilde{\mathcal{Q}}$ and also obtain its latent representation $\widetilde{\mathcal{I}}$ through VAE’s encoder. The spatial resolution of $\widetilde{\mathcal{Q}}$ is pre-upscaled by several auxiliary convolutional layers. Besides, we employ CLIP’s image encoder to extract the vectorized embedding $\tilde{\mathbf{c}}$ for cross-attention within U-Net’s transformer blocks. For the rendered depth map $\mathcal{D}$, we utilize its latent representation $\widetilde{\mathcal{D}}$ as the geometric prior. Overall, we concatenate $\widetilde{\mathcal{Q}}$, $\widetilde{\mathcal{I}}$, and $\widetilde{\mathcal{D}}$ to deduce the resulting conditioning signal $\widetilde{\mathcal{S}}$.
@@
Empirically, we introduce another parameterization form of v-prediction [86] for training the denoising U-Net, whose input channels and cross-attention dimensions are adapted to match with our prior-conditioned latent representations and the CLIP-encoded image embeddings.
III-C Noise-Resistant Inverse Normal Rendering
Figure 3: Results of applying our proposed SuperCarver for texture-consistent 3D geometry super-resolution.
@@ III-C Noise-Resistant Inverse Normal Rendering
When finishing the training of our normal diffusion model, we can render multi-view 2D images from different viewpoints $\{\mathbf{v}_{k}\}_{k=1}^{K}$, and perform inference to correspondingly predict a collection of detail-rich normal maps $\{\mathcal{H}_{k}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{K}$ to facilitate multi-view inverse rendering.
@@
To transfer the underlying 3D structural patterns implied by the predicted 2D normal maps to the input mesh surface, the most common practice is to perform gradient-based mesh refinement through inverse rendering. However, it is observed that updating the mesh geometry from potentially imperfect normal map predictions is prone to causing obvious surface degradations, even with adaptations of various existing state-of-the-art (either explicit [68] or implicit [80]) mesh optimization approaches. Hence, we are strongly motivated to develop a highly noise-resistant scheme better suited to our setting of supplementing new details to a given coarse surface in a non-destructive manner.
@@
We introduce a simple yet effective distance field deformation approach, which works in a “value-fixing, grid-moving” manner. Specifically, assuming that the original mesh $\mathcal{M}=(\mathcal{V},\mathcal{F})$ has already been normalized into the unit sphere, we start by constructing uniform 3D grid points $\mathcal{G}\in\mathbb{R}^{r\times r\times r\times 3}$ within $[-1,1]^{3}$ and computing per-grid signed distance values $\mathcal{X}\in\mathbb{R}^{r\times r\times r}$. Different from common ways of implementing deformable distance field representations via only updating $\mathcal{X}$ or jointly updating both $\mathcal{G}$ and $\mathcal{X}$, we choose a subtle strategy that fixes the distance values $\mathcal{X}$ while separately learning zero-initialized offsets $\mathcal{O}\in\mathbb{R}^{r\times r\times r\times 3}$ to move initial grid positions $\mathcal{G}$. In comparison, standard distance field refinement requires higher resolutions of fixed uniform grids; otherwise, detailed structures cannot be well captured even with accurate distance values. Additionally, fixing pre-computed distance values hinders the optimization process from drastically changing the original mesh structure even with obviously erroneous normal prediction in certain areas.
@@
To achieve this, we employ a differentiable implementation of the classic iso-surface extraction algorithm of DMC [73], which we denote as $\Psi(\cdot;\cdot)$, for producing the updated mesh $\mathcal{M}^{u}=(\mathcal{V}^{u},\mathcal{F}^{u})$. Then, a differentiable rasterizer $\mathcal{R}(\cdot;\cdot)$ is integrated to render the correspondingly updated normal maps $\{\mathcal{H}_{k}^{u}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{K}$ with respect to each viewpoint $\mathbf{v}_{k}$. The whole processing procedure can be formulated as:
@@
where the first term aligns all pairs of the diffusion-predicted and mesh-rendered normal maps, the last two terms denote mesh vertex and normal regularizations weighted by factors $\omega_{s}$ and $\omega_{n}$. For each vertex $\mathbf{p}_{i}$ in the updated mesh and its neighboring vertices $\Omega_{i}=\{\mathbf{p}_{i,j}\}$, the Laplacian smoothing is computed by:
@@
After iterative optimization, we moderately apply the Taubin smoothing [87] as a post-processing step to mitigate meaningless surface wrinkles.
IV Experiments
This section begins by introducing our main technical implementations in terms of data preparation and framework hyperparameters. For performance evaluations, we qualitatively illustrate the boosting effects of applying the proposed SuperCarver to a collection of diverse real-world 3D assets as well as synthesized mesh models from different 3D AIGC methods. To validate the superiority of our core technical components, we make comparisons with state-of-the-art normal prediction and mesh optimization approaches. We further conduct a series of ablation studies and analyses. In the end, we also discuss our current limitations and potential future explorations.
Figure 5: Results of normal map predictions generated from (c) GeoWizard [24], (d) StableNormal [88], and (e) Ours, where (a) and (b) show the corresponding color images and ground-truth high-fidelity normal maps.
Figure 6: Rendering effects of (c) our optimized mesh and (d) attaching the bump map obtained from PBRBoost [68], given the input textured coarse mesh in (a) and (b).
Training Dataset Preparation. We used the publicly available Objaverse [6] dataset comprising 800K+ 3D assets to produce our training data. Through a series of meticulously designed filtering protocols and cleaning-up procedures, we ultimately acquired around 80K high-quality mesh models.
To create the paired low-poly and high-poly normal renderings, we performed geometry simplification on each high-poly training mesh model based on manifold reconstruction [90], and further removed detailed structures using [87]. We further manually annotated a smaller portion of unqualified training samples, which were used for training a separate neural model to purify the whole training set.
Implementation Details. We employed the popular diffusion architecture of Stable Diffusion V2.1 [5] and performed fine-tuning on 16 NVIDIA Tesla V100 GPUs for 40K steps with a total batch size of 128. We applied the AdamW [91] optimizer with cosine annealing, reducing the learning rate from 5e-5 to 1e-5. For both training and inference, the image resolution is uniformly configured to 512$\times$512.
For iterative mesh optimization, we adopted Nvdiffrast [92] for efficient differentiable rendering from 12 camera poses (8 evenly-spaced horizontal views with azimuth angles from 0 to 360 and 4 downward-looking (elevated by 30) views with azimuth angles [45, 135, 225, 315]). Camera intrinsics were kept the same as our configurations in BlenderProc. The hyperparameters were empirically tuned as $r=512$, $\tau=0.5$, $\omega_{s}=0.25$, $\omega_{n}=0.01$ for 200 iterations.
When deployed on NVIDIA L20 GPU (mid-range product for AI inference), the average time consumption for the whole geometry super-resolution workflow is around 1$\sim$2 minutes, which is thought to be adequately efficient.
IV-A Performance Evaluation
Figure 7: Typical surface damages produced by FlexiCubes [80].
Considering the lack of specialized researches on our targeted task of texture-consistent geometry super-resolution, we provided qualitative visualizations for the outputs of applying SuperCarver to diverse geometrically-coarse textured meshes. As illustrated in Figure 3, our approach is capable of supplementing realistic and expressive surface details under a variety of testing cases.
In addition to real-world 3D assets, we further experimented with AI-generated mesh assets by applying SuperCarver to the textured mesh models generated from advanced 3D generation frameworks. We adopted InstantMesh [9], Unique3D [12], and CLAY [13], covering representative generation paradigms of feed-forward large reconstruction models, multi-view diffusion models, and native 3D latent diffusion models. As compared in Figure 4, various input meshes are effectively boosted by our approach to produce realistic and expressive 3D surface details well aligned with the corresponding texture appearances.
Comparison on Normal Prediction. To evaluate the effectiveness of our specialized normal diffusion architecture, we made comparisons with two advanced normal diffusion frameworks, including GeoWizard [24] and StableNormal [88]. To facilitate quantitative evaluations, we manually selected 50 detail-rich textured meshes as ground-truths from the widely-used Objaverse [6] and ABO [93] shape repositories. We applied surface simplification and smoothing to produce geometrically-coarse meshes as the input samples.
As compared in Figure 5, our normal map predictions show better details and avoid obviously erroneous areas. Besides, we further deduced the metric of mean angular error (MAE) between ground-truth and predicted normal maps, where we masked invalid backgrounds for the outputs of GeoWizard and StableNormal. As shown in Table I, thanks to our specialized diffusion architecture, our normal map predictions significantly outperform the competing approaches.
TABLE I: Quantitative comparison of normal map prediction accuracy under the metric of mean angular error (MAE).
@@ IV-A Performance Evaluation
TABLE II: Normal map diffusion accuracy of different baseline schemes. “Stoc.” denotes the straightforward implementation of stochastic diffusion, then “w/o Sem.” and “w/o Geo.” respectively represent removing the high-level semantics and low-level depth priors.
@@
TABLE III: Quantitative comparison of our different baseline design choices for achieving distance field deformation. “Only Grids” represents the original design, “Only Dist.” and “Joint G&D” are the corresponding two variants.
@@
TABLE IV: Quantitative evaluation of cross-view consistency under the metric of mean angular error (MAE).
Overall Normal Diffusion Error Cross-View Normal Inconsistency
9.05 7.24
In addition, we also included PBRBoost [25] (an advanced material generation framework), since one of its outputs is the 2D bump maps associated with the base meshes, enabling us to compare its rendering effects when attaching the bump maps as height information. As shown in Figure 6, the actual fidelity of the enhanced geometric cues from PBRBoost is still far from satisfactory.
Comparison on Mesh Optimization. To facilitate 3D surface geometry updating, many recent mesh generation frameworks [94, 95, 96, 12] introduce the CR [68] algorithm with joint optimization and remeshing, which is regarded as the most relevant competitor. In fact, we had previously attempted to adapt FlexiCubes [80], an advanced iso-surface representation, to achieve gradient-based geometry refinement starting from the initially given mesh surface. However, through extensive attempts, we empirically found that it easily leads to severe surface damages, as illustrated in Figure 7. We reason that this is attributable to its significantly increased deformation flexibility (which should have been its major strength) amplifies the impact of erroneous normal pixels. As shown in Figure 8, for typical hard cases with less accurate normal predictions, CR tends to produce noisy and degraded surfaces.
IV-B Ablation Study
To evaluate the effectiveness and necessity of the core components in our entire workflow, we conducted ablation study with different baseline schemes. We began with a straightforward baseline normal diffusion framework as formulated in Eq. (1) and Eq. (2). Besides, since prior guidance is also a critical condition for generating robust and accurate normal maps, we can design another two variants via removing the semantic cues extracted from DINO features or the geometric cues extracted from depth maps.
As reported in Table II, directly fine-tuning a standard diffusion model conditioned on color images lead to significantly degraded performance, with the MAE metric of 21.64. More importantly, as illustrated in Figure 9, the stochastic nature inevitably results in unstable predictions. As reported in the last two columns of Table II, the joint injection of semantic and geometric priors are indeed beneficial.
For inverse normal rendering, we experimented with another two distance field deformation schemes, i.e., fixing grids while optimizing distances; jointly updating both grids and distances. To obtain quantitative results, we computed the MAE metric of normal maps rendered from optimized meshes (note that Table I and Table II are targeted at diffused normal maps). As evidenced in Table III, the two variants show sub-optimal performances, in that only optimizing distances usually causes weakened details while the joint optimization strategy easily suffers from surface distortion especially when the diffused normal maps contain relatively larger errors.
Figure 10: Illustration of our multi-view normal predictions.
Additionally, we quantitatively measured normal deviations across views in Table IV, in which the average MAE is 7.24. Comparing this metric with the overall normal prediction error with respect to ground-truth normal maps and the overall performances of other state-of-the-art competitors as reported in Table I, we can observe that our view consistency performance is satisfactory enough, as visualized in Figure 10.
IV-C Limitations & Future Works
Figure 11: Illustration of over-detailization in Hunyuan3D [97].
Figure 12: Illustration of applying SuperCarver to the generated textured meshes of Hunyuan3D [97] (left) to further enhance surface detail quality (right).
As the first learning-based solution for texture-consistent 3D geometry super-resolution, SuperCarver demonstrates satisfactory fidelity and applicability. Nevertheless, as a preliminary attempt in this direction, our current technical implementation still has several aspects of limitations for future explorations.
With finite image resolution, our normal boosting component cannot effectively enhance the detail patterns of elongated objects (e.g., knives, swords), since the rendered areas are too small to provide adequate appearance information. Localized processing would address the resolution issue.
Hindered by the inconsistent quality of training data, over-detailization occurs in certain cases. Theoretically, identifying geometry-irrelevant textures is supposed to be unified within the overall learning process when training with large volumes of data. Therefore, more stringent data selection protocols are needed to ensure geometry-texture consistency in the training samples. Still, it is worth noting that texture overfitting remains unresolved in the current 3D generation community. As shown in Figure 11, even the latest state-of-the-art foundation model Hunyuan3D [97] still suffers from the occurrence of overly detailed 3D geometric structures. (Note that SuperCarver works consistently well on Hunyuan3D-generated assets as illustrated in Figure 12.)
Additionally, the inherent properties of the adopted signed distance field representation restrict our capabilities to handle open surfaces and highly complicated topologies. Hence, we need to explore differentiable iso-surface extractors built upon more generic implicit geometry representations.
Notably, our second stage of mesh detail generation relies on iterative gradient-based optimization. One promising direction is to build feed-forward neural architectures to further improve the geometric quality and running efficiency.
V Conclusion